OpenStudy (anonymous):

Write the equation of the line containing the points (-1,3) and (2,4) in y=mx+b form. Help!

jimthompson5910 (jim_thompson5910):

First find the slope

jimthompson5910 (jim_thompson5910):

Any ideas on how to do that?

OpenStudy (noseboy908):

rise/run. aka: Y2-Y1/X2-X1

OpenStudy (anonymous):

\[\frac{ 4-3 }{2+1}\]

OpenStudy (anonymous):

Yes, the slope formula. m= y2-y1 over x2-x1

jimthompson5910 (jim_thompson5910):

good, so the slope is ???

OpenStudy (anonymous):

find the slope with that then chose one set of coordinates to replace the x- and y- values.

OpenStudy (anonymous):

1/3 ?

jimthompson5910 (jim_thompson5910):

yes, good

jimthompson5910 (jim_thompson5910):

so m = 1/3 now plug in any of the two points, so if you pick the first point, then you plug in x = -1 and y = 3 into y = mx+b to get y = mx+b y = 1/3x + b 3 = (1/3)(-1) + b Keep going to solve for b

OpenStudy (anonymous):

Thank you guys!! That was helpful @a0zzy and @jim_thompson5910

OpenStudy (anonymous):

no problem

jimthompson5910 (jim_thompson5910):

you're welcome